OpenStudy (anonymous):

Hello. I am Stuck. All help is appreciated. Question: Expand - (5x^2-6y^2)^2 . Many Thanks.

OpenStudy (compassionate):

- (5x^2-6y^2)^2 - ( 5 x ^2 -6 y^2 ) ^2 - ( 5 x^2 -6 y^2 ) ^2 - ( 5 x^2 -6 y^2 )^2

OpenStudy (compassionate):

But, is the "-" in front a negative sign, or is that just a comma?

OpenStudy (sidsiddhartha):

lol

OpenStudy (sidsiddhartha):

i think hshould use (a-b)^2=a^2-2ab+b^2

OpenStudy (anonymous):

( five x squared minus six y squared ) squared

OpenStudy (sidsiddhartha):

*it will be (5x^2-6y^2)^2

OpenStudy (sidsiddhartha):

\[(5x^2-6y^2)^2=(5x^2)^2-2.5x^2.6y^2+(6y)^2\]

OpenStudy (anonymous):

\[where does the -2.5x^{2} come from\]

OpenStudy (sidsiddhartha):

\[used\\(a-b)^2=a^2-2.a.b+b^2\]

OpenStudy (sidsiddhartha):

here (.) means multiplication

OpenStudy (sidsiddhartha):

so finally\[(5x^2-6y^2)=25x^4-60x^2y^2+36y^2\]

OpenStudy (anonymous):

why dosen't Y get a 4 as an exponent ?

OpenStudy (sidsiddhartha):

yeah my bad it will get a 4 exponent :)

OpenStudy (anonymous):

Thank you so much. How do I give you a medal?

OpenStudy (sidsiddhartha):

just click "best response" button :)

OpenStudy (anonymous):

I got it. Thanks again. I need to practice more.

OpenStudy (sidsiddhartha):

yw!!and good luck :)